# Disentangling aleatoric and epistemic risks in bus fleet control: A stable and robust ensemble DRL approach [^1] Code available at https://github.com/erzhu419/RE-SAC

### 19 Mar 2026 v1

@@ Abstract
Empirically, experiments in a realistic bidirectional bus corridor simulation demonstrate that RE-SAC achieves the highest and most stable cumulative reward (approximately $0.4\times 10^{6}$) among all baselines, compared to $0.55\times 10^{6}$ for vanilla SAC and $1.2\times 10^{6}$ for the epistemic-only ablation. The Mahalanobis rareness analysis further confirms that RE-SAC reduces Oracle Q-value estimation error by up to $62\%$ in rare out-of-distribution states relative to SAC (average Oracle MAE of 1647 vs. 4343), demonstrating superior worst-case robustness under high traffic variability.
@@ 1 Introduction
Building upon our previous work on single-agent robust control for bus fleets [38], this paper introduces the RE-SAC framework. Our contributions are summarized as follows:
@@ 4.1 Theoretical formulation: Robustness via regularization
This formulation allows us to bypass the complexity of finding the worst-case distribution $p$ by directly penalizing the function space complexity, akin to the equivalence between robustness and regularization shown in support vector machines [34] (see Appendix B). When $\mathcal{F}$ is the set of 1-Lipschitz functions, $d_{\mathcal{F}}$ coincides with the 1-Wasserstein distance.
@@
where $\text{Lip}(V_{\theta})$ is the Lipschitz constant of the value network. Since the Lipschitz constant of a deep network is upper-bounded by the product of the spectral norms of its weight matrices, i.e., $\text{Lip}(V_{\theta})\leq\prod_{l}\|W_{l}\|_{2}$, robustness against aleatoric perturbations can be achieved by regularizing the critic network weights. In practice, we use the $\ell_{1}$-norm as a computationally efficient surrogate. For the $L_{\infty}$ perturbation model relevant to bounded traffic states, the per-layer sensitivity satisfies $\|W_{l}\delta\|_{\infty}\leq\|W_{l}\|_{\infty\to\infty}\|\delta\|_{\infty}\leq\|\text{vec}(W_{l})\|_{1}\|\delta\|_{\infty}$, so penalizing $\sum_{l}\|W_{l}\|_{1}$ directly controls the worst-case $L_{\infty}$ sensitivity of the network (see Appendix D.5).
@@ 4.2 Disentangling epistemic and aleatoric risks
2. Aleatoric risk ($\Gamma_{ale}$): This term accounts for the irreducible noise in the environment (e.g., stochastic passenger arrivals). Based on the IPM duality proven in [39], this risk is controlled by the Lipschitz constant of the value function: $\Gamma_{ale}(V)=\varepsilon\cdot\mathrm{Lip}(V_{\theta})$, where $V_{\theta}$ is parameterized by the current critic weights $\theta$. A non-smooth $V$ with large Lipschitz constants is highly sensitive to aleatoric perturbations. As established in Eq. (5), bounding $\text{Lip}(V_{\theta})$ via weight norm regularization ($\|W\|_{1}$) provides structural robustness to inherent noise. Crucially, during each Bellman backup the weights $\theta$ are held fixed (updated only afterward via gradient descent), so the aleatoric penalty collapses to a fixed scalar $\kappa=\lambda_{ale}\sum_{l}\|W_{l}^{(\theta)}\|_{1}$ that is independent of the Q-function being evaluated (see Appendix D.5).
According to the equivalence theory of Osogami [25] (detailed in Appendix B), this additive decomposition corresponds to an uncertainty set $\mathcal{P}_{s,a}$ that is the intersection of a data-density-based set and a function-class-based IPM set. This disentanglement allows the agent to maintain high $Q$-values in noisy but well-explored states, thereby avoiding the underestimation trap while maintaining structural robustness.
@@ 4.4 Epistemic risk mitigation via $Q$-ensemble
where $\bar{Q}(s,a)$ is the mean prediction of the ensemble.
@@
$\mathcal{T}_{epi}Q(s,a)=R(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim p^{\circ}}[\bar{Q}(s^{\prime},a^{\prime})-\lambda_{epi}\Gamma_{epi}(s^{\prime},a^{\prime})].$ (10)
@@
Synergy between disentangled uncertainties: The synergy between the $Q$-ensemble and IPM regularization is the key to preventing $Q$-value poisoning. As shown in our experiments, using the ensemble alone often fails in high-variance transit environments because it treats aleatoric noise as epistemic gaps, leading to excessive pessimism in well-explored but noisy states. Conversely, using only weight regularization fails to detect OOD regions. By disentangling these two, the ensemble captures the data-density-based risk, while the IPM regularization handles the local-perturbation-based risk. This combined defense ensures that the $Q$-value remains stable and accurately reflects the worst-case scenario without falling into the underestimation trap identified by [20].
@@ 4.5.1 Critic objective with disentangled penalties
$y=r+\gamma\left(\mathbb{E}_{a^{\prime}\sim\pi_{\theta}}\left[\frac{1}{K}\sum_{k=1}^{K}\hat{Q}_{\phi_{k}}(s^{\prime},a^{\prime})-\lambda_{epi}\cdot\Gamma_{epi}(s^{\prime},a^{\prime})-\lambda_{ale}\sum_{l=1}^{L}\|W_{l}^{(\theta)}\|_{1}\right]-\alpha\log\pi_{\theta}(a^{\prime}|s^{\prime})\right),$ (11)
where $\hat{Q}_{\phi_{k}}$ denotes the target $Q$-networks, $\Gamma_{epi}(s^{\prime},a^{\prime})$ is the epistemic penalty (ensemble variance, Eq. (LABEL:eq:gamma_epi)), and $\lambda_{ale}\sum_{l}\|W_{l}^{(\theta)}\|_{1}$ is the aleatoric scalar $\kappa$ computed from the frozen target-network weights $\theta$ (cf. §4.2 and Appendix D.5). Since $\kappa$ does not depend on which Q-function we evaluate, it cancels in the contraction proof.
@@
where $\|W_{l}^{(k)}\|_{1}$ is the $\ell_{1}$-norm of the weights in the $l$-th layer, ensuring Lipschitz continuity (see Appendix D.5 for a contraction proof and Appendix D for a theoretical comparison with SAC’s robustness). The term $\sigma_{ens}(s,a)=\text{Std}(\{Q_{\phi_{k}}(s,a)\}_{k=1}^{K})$ penalizes large ensemble disagreement during training.
@@ 4.5.3 Algorithm pseudocode
   $y=r_{t}+\gamma\left(\text{Mean}(Q_{\phi^{\prime}_{k}}(s_{t+1},a^{\prime}))-\lambda_{epi}\Gamma_{epi}-\lambda_{ale}\sum_{l}\|W_{l}\|_{1}-\alpha\log\pi_{\theta}(a^{\prime}|s_{t+1})\right)$.
@@ 6 Conclusion
This paper presents the RE-SAC framework, a robust single-agent reinforcement learning approach for bus fleet holding control that explicitly disentangles aleatoric and epistemic uncertainties. Grounded in the theoretical equivalence between risk-sensitivity and robustness established by Osogami [25], our method employs two complementary mechanisms: IPM-based weight regularization to enforce Lipschitz continuity of the value function against inherent environmental stochasticity, and a diversified $Q$-ensemble to penalize overconfident estimates in data-sparse regions of the state-action space.
Our theoretical analysis demonstrates that this disentangled approach retrieves polynomial sample complexity bounds, bypassing the exponential barrier inherent in direct risk-sensitive optimization. Empirically, experiments on a realistic bidirectional bus simulation confirm that RE-SAC achieves the highest and most stable cumulative reward among all baselines, including vanilla SAC, DSAC, and BAC. The ablation studies reveal that relying solely on ensemble variance (Epistemic-Only) leads to catastrophic value collapse in high-variance transit environments, while using only weight regularization (Aleatoric-Only) fails to detect out-of-distribution states. The Mahalanobis rareness analysis further confirms that RE-SAC maintains accurate $Q$-value estimates even under highly rare traffic states, exhibiting superior worst-case robustness.
@@ A.1 Risk-Sensitive RL: The exponential barrier
A.2 RE-SAC: Retrieving polynomial bounds
Why does RE-SAC achieve polynomial complexity while the equivalent Risk-Sensitive problem is exponential? The key lies in the surrogate objective.
Directly Optimizing $\mathbb{E}[\exp(\beta R)]$ (Risk-Sensitive) requires accurate estimation of the tail distribution, which is statistically expensive (hence $\exp(H)$). However, RE-SAC leverages the specific duality shown by Xu et al. (Appendix B), which maps a specific form of robustness (IPM-based) to Norm Regularization.
@@ A.2 RE-SAC: Retrieving polynomial bounds
This transformation is crucial: it converts the "Hard" Risk-Sensitive problem into a "Standard" Risk-Neutral MDP problem where the reward is simply augmented by a deterministic regularizer $\tilde{R}(s,a)=R(s,a)-\lambda\|W\|^{2}$. Since solving a standard MDP with bounded rewards has polynomial regret $\tilde{O}(\sqrt{H^{3}S^{2}AT})$, RE-SAC inherits this efficiency. We effectively bypass the need to statistically estimate the heavy-tailed distribution (which requires exponential samples) by enforcing a structural constraint (smoothness) that serves as an efficient proxy for risk aversion.
$\text{Regret}_{RE-SAC}(T)\leq\tilde{O}(\sqrt{H^{3}S^{2}AT}).$ (20)
Remark. This analysis applies to the tabular setting with finite $S,A$. The practical claim is that our regularization-based approach avoids the exponential dependence on the risk parameter $\beta$ that is inherent to direct risk-sensitive optimization, motivating it as a tractable surrogate in the deep RL setting.
Corollary (Sample complexity): Similar to the risk-sensitive case, we derive the sample complexity for RE-SAC. Proof Sketch: To ensure the average optimality gap $\Delta(T)\leq\epsilon$:
$\frac{\tilde{O}(\sqrt{H^{3}S^{2}AT})}{T}$ $\leq\epsilon$
$\tilde{O}\left(\frac{\sqrt{H^{3}S^{2}A}}{\sqrt{T}}\right)$ $\leq\epsilon$
$T$ $\geq\tilde{O}\left(\frac{H^{3}S^{2}A}{\epsilon^{2}}\right).$ (21)
Appendix B The equivalence chain: From risk to regularization

### 21 Mar 2026 v2

# RE-SAC: Disentangling aleatoric and epistemic risks in bus fleet control: A stable and robust ensemble DRL approach [^1] Code available at https://github.com/erzhu419/RE-SAC

@@ Abstract
Empirically, experiments in a realistic bidirectional bus corridor simulation demonstrate that RE-SAC achieves the highest and most stable cumulative reward (approximately $0.4\times 10^{6}$) among all baselines, compared to $0.55\times 10^{6}$ for vanilla SAC and $1.2\times 10^{6}$ for the epistemic-only ablation. The Mahalanobis rareness analysis further confirms that RE-SAC reduces Oracle Q-value estimation error by up to $62\%$ in rare out-of-distribution states relative to SAC (average Oracle MAE of 1647 vs. 4343), demonstrating superior worst-case robustness under high traffic variability. Our theoretical analysis further reveals that the frozen-parameter design (target networks, held-fixed critic weights) is not merely a practical stabilisation heuristic but a structurally necessary condition for Bellman operator contractiveness under multi-penalty regularisation—an insight with implications beyond transit control.
@@ 1 Introduction
Definition 1.1 ($Q$-value Poisoning).
Let $Q^{*}$ be the optimal $Q$-function and $Q_{t}$ be the learned $Q$-function after $t$ gradient steps. Let $\sigma^{2}(s,a):=\mathrm{Var}_{P}[V_{t}(s^{\prime})\mid s,a]$ denote the local aleatoric variance at $(s,a)$. We say the learning process suffers from $Q$-value poisoning if there exist a threshold $\sigma^{2}_{0}>0$ and a non-vanishing set of state-action pairs $\mathcal{P}_{t}\subseteq\mathcal{S}\times\mathcal{A}$ such that:
  1. (i)
    Selective underestimation in high-variance regions: $Q_{t}(s,a)<Q^{*}(s,a)$ for all $(s,a)\in\mathcal{P}_{t}$ with $\sigma^{2}(s,a)\geq\sigma^{2}_{0}$; and
  2. (ii)
    Accurate estimation in low-variance regions: $|Q_{t}(s,a)-Q^{*}(s,a)|\leq\epsilon$ for $(s,a)\notin\mathcal{P}_{t}$ with $\sigma^{2}(s,a)<\sigma^{2}_{0}$.
This pattern—accurate in low-noise states, poisoned in high-noise states—is structurally distinct from classical overestimation bias ($Q_{t}\geq Q^{*}$ uniformly) and from the Deadly Triad (divergence of $\|Q_{t}\|_{\infty}\to\infty$); see Appendix H.
@@
Building upon our previous work on single-agent robust control for bus fleets [20], this paper introduces the RE-SAC framework. More broadly, our results suggest that contraction in regularised reinforcement learning is not a generic property of penalised Bellman operators, but depends critically on whether the penalty term is fixed or value-function-dependent during each backup step. Making this structural assumption explicit—and verifying it formally—is essential for a correct theoretical understanding of modern deep RL algorithms. Our contributions are summarized as follows:
@@
  • Structural stability characterisation: We identify a sharp boundary for contraction in multi-penalty Bellman operators: penalties that are fixed per backup step (frozen critic weights, target networks) preserve $\gamma$-contraction, while Q-dependent penalties provably break it. Both the positive and negative results are machine-verified in Lean 4 with no sorry (Proof.lean, Counterproof.lean). This reveals a necessary design condition for stability in regularised RL—the frozen-parameter architecture used in practice (DQN, SAC, TD3) is not merely convenient but structurally required.
@@ 4.1 Theoretical formulation: Robustness via regularization
This formulation allows us to bypass the complexity of finding the worst-case distribution $p$ by directly penalizing the function space complexity, akin to the correspondence between robustness and regularization shown in support vector machines [15] (see Appendix B). When $\mathcal{F}$ is the set of 1-Lipschitz functions, $d_{\mathcal{F}}$ coincides with the 1-Wasserstein distance.
@@
where $\text{Lip}(V_{\theta})$ is the Lipschitz constant of the value network. Since the Lipschitz constant of a deep network is upper-bounded by the product of the spectral norms of its weight matrices, i.e., $\text{Lip}(V_{\theta})\leq\prod_{l}\|W_{l}\|_{2}$, robustness against aleatoric perturbations can be achieved by regularizing the critic network weights. In practice, we use the $\ell_{1}$-norm as a computationally efficient surrogate. For the $L_{\infty}$ perturbation model relevant to bounded traffic states, the per-layer sensitivity satisfies $\|W_{l}\delta\|_{\infty}\leq\|W_{l}\|_{\infty\to\infty}\|\delta\|_{\infty}\leq\|\text{vec}(W_{l})\|_{1}\|\delta\|_{\infty}$, so penalizing $\sum_{l}\|W_{l}\|_{1}$ directly controls the worst-case $L_{\infty}$ sensitivity of the network (see Appendix E for a machine-verified contraction proof in finite state-action spaces, with the Lean 4 files Proof.lean establishing contractivity and Counterproof.lean establishing that Q-dependent penalties strictly break it, plus a continuous-space extension via the Banach Fixed-Point Theorem).
@@ 4.2 Disentangling epistemic and aleatoric risks
2. Aleatoric risk ($\Gamma_{ale}$): This term accounts for the irreducible noise in the environment (e.g., stochastic passenger arrivals). Based on the IPM duality proven in [16], this risk is controlled by the Lipschitz constant of the value function: $\Gamma_{ale}(V)=\varepsilon\cdot\mathrm{Lip}(V_{\theta})$, where $V_{\theta}$ is parameterized by the current critic weights $\theta$. A non-smooth $V$ with large Lipschitz constants is highly sensitive to aleatoric perturbations. As established in Eq. (5), bounding $\text{Lip}(V_{\theta})$ via weight norm regularization ($\|W\|_{1}$) provides structural robustness to inherent noise. Crucially, during each Bellman backup the weights $\theta$ are held fixed (updated only afterward via gradient descent), so the aleatoric penalty collapses to a fixed scalar $\kappa=\lambda_{ale}\sum_{l}\|W_{l}^{(\theta)}\|_{1}$ that is independent of the Q-function being evaluated (see Appendix E).
According to the correspondence theory of Osogami [13] (detailed in Appendix B), this additive decomposition corresponds to an uncertainty set $\mathcal{P}_{s,a}$ that is the intersection of a data-density-based set and a function-class-based IPM set. This disentanglement allows the agent to maintain high $Q$-values in noisy but well-explored states, thereby avoiding the underestimation trap while maintaining structural robustness. A formal justification that the two penalties target genuinely distinct sources of uncertainty—including a proof that $\Gamma_{epi}\to 0$ as data coverage grows while $\kappa$ remains bounded away from zero—is provided in Appendix F.
@@ 4.4 Epistemic risk mitigation via $Q$-ensemble
where $\bar{Q}(s,a)$ is the mean prediction of the ensemble. Crucially, $\Gamma_{epi}$ is computed from the frozen target networks $\hat{Q}_{\phi^{\prime}_{k}}$, not the online networks being updated. It is therefore a fixed function of $(s,a)$ during each Bellman backup, independent of the Q-function being evaluated—exactly the structural property required for contraction (Appendix E.1, S3).
@@
$\mathcal{T}_{epi}Q(s,a)=R(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim p^{\circ}}[\bar{Q}(s^{\prime},a^{\prime})]-\gamma\,\lambda_{epi}\Gamma_{epi}(s,a).$ (10)
@@
Synergy between disentangled uncertainties: The synergy between the $Q$-ensemble and IPM regularization is the key to preventing $Q$-value poisoning. As shown in our experiments, using the ensemble alone often fails in high-variance transit environments because it treats aleatoric noise as epistemic gaps, leading to excessive pessimism in well-explored but noisy states. Conversely, using only weight regularization fails to detect OOD regions. By disentangling these two, the ensemble captures the data-density-based risk, while the IPM regularization handles the local-perturbation-based risk. This combined defense ensures that the $Q$-value remains stable and accurately reflects the worst-case scenario without falling into the underestimation trap identified by [4]. A formal condition on the penalty coefficients ($\lambda_{ale}$, $\lambda_{epi}$) guaranteeing that the fixed point $Q^{*}_{robust}$ remains strictly positive—and thus the policy avoids degeneration—is given in Appendix G.
@@ 4.5.1 Critic objective with disentangled penalties
$y=r+\gamma\left(\mathbb{E}_{a^{\prime}\sim\pi_{\theta}}\left[\frac{1}{K}\sum_{k=1}^{K}\hat{Q}_{\phi_{k}}(s^{\prime},a^{\prime})-\lambda_{epi}\cdot\Gamma_{epi}(s^{\prime},a^{\prime})\right]-\alpha\log\pi_{\theta}(a^{\prime}|s^{\prime})\right),$ (11)
where $\hat{Q}_{\phi_{k}}$ denotes the target $Q$-networks and $\Gamma_{epi}(s^{\prime},a^{\prime})$ is the epistemic penalty (ensemble variance, Eq. (9)). Aleatoric risk is mitigated not by subtracting $\kappa$ from the Bellman target—which would push already-negative bus-holding rewards further negative, violating the feasibility condition of Appendix G—but by penalising large weight norms in the critic loss $\mathcal{L}_{Q}$ (see below). [^3] Concretely, the implementation adds $\lambda_{ale}\kappa$ to the target Q-values (equivalent to a signed intrinsic reward that incentivises Lipschitz-smooth critics), and also adds $\lambda_{ale}\kappa$ in the actor’s LCB objective, ensuring joint consistency. The contraction proof is unaffected: $\kappa$ is still a fixed scalar per Bellman backup.
@@
where $\|W_{l}^{(k)}\|_{1}$ is the $\ell_{1}$-norm of the weights in the $l$-th layer, ensuring Lipschitz continuity (see Appendix E for a contraction proof and Appendix D for a theoretical comparison with SAC’s robustness). The term $\sigma_{ens}(s,a)=\text{Std}(\{Q_{\phi_{k}}(s,a)\}_{k=1}^{K})$ penalizes large ensemble disagreement during training.
Remark 4.1 (Scope of the contraction guarantee).
The contraction result applies to the operator under the frozen-parameter interpretation (target networks for $\Gamma_{epi}$, held-fixed critic weights for $\kappa$) and does not imply convergence of the full learning dynamics. Contraction is preserved when penalties are fixed per backup step; it provably fails when penalties depend on the value function being updated (demonstrated in Counterproof.lean). We use an upper bound on the aleatoric penalty as surrogate (Appendix E.1, S2): this is conservative but does not affect the existence or uniqueness of the fixed point.
Remark 4.2 (Scope of the contraction guarantee).
The contraction result applies to the operator under the frozen-parameter interpretation (target networks for $\Gamma_{epi}$, held-fixed critic weights for $\kappa$) and does not imply convergence of the full learning dynamics. Contraction is preserved when penalties are fixed per backup; it may fail when penalties depend on the value function being updated—an impossibility demonstrated in Counterproof.lean. We analyze a surrogate operator using an upper bound on the original aleatoric penalty (cf. Appendix E.1, S2); this is conservative but does not affect the existence or uniqueness of the fixed point.
@@ 4.5.3 Algorithm pseudocode
   $y=r_{t}+\gamma\left(\text{Mean}(Q_{\phi^{\prime}_{k}}(s_{t+1},a^{\prime}))-\lambda_{epi}\Gamma_{epi}-\alpha\log\pi_{\theta}(a^{\prime}|s_{t+1})\right)$. // aleatoric reg. in $\mathcal{L}_{Q}$ below
@@ 6 Conclusion
This paper presents the RE-SAC framework, a robust single-agent reinforcement learning approach for bus fleet holding control that explicitly disentangles aleatoric and epistemic uncertainties. Grounded in the theoretical correspondence between risk-sensitivity and robustness established by Osogami [13], our method employs two complementary mechanisms: IPM-based weight regularization to enforce Lipschitz continuity of the value function against inherent environmental stochasticity, and a diversified $Q$-ensemble to penalize overconfident estimates in data-sparse regions of the state-action space.
Our theoretical analysis identifies a sharp structural boundary for contraction in multi-penalty Bellman operators: the frozen-parameter design (target networks for epistemic variance, held-fixed critic weights for aleatoric penalty) is both sufficient for $\gamma$-contraction and necessary—Q-dependent penalties provably break it. Both results are machine-verified in Lean 4 with no sorry, and a verifiable feasibility condition on the penalty coefficients guarantees policy non-degeneration (Appendix G). This structural insight—that the standard deep RL architecture of frozen target networks is not merely a stabilisation heuristic but a required condition for operator contractiveness—may inform the design of future multi-penalty RL algorithms beyond the transit domain. Empirically, experiments on a realistic bidirectional bus simulation confirm that RE-SAC achieves the highest and most stable cumulative reward among all baselines, including vanilla SAC, DSAC, and BAC. The ablation studies reveal that relying solely on ensemble variance (Epistemic-Only) leads to catastrophic value collapse in high-variance transit environments, while using only weight regularization (Aleatoric-Only) fails to detect out-of-distribution states. The Mahalanobis rareness analysis further confirms that RE-SAC maintains accurate $Q$-value estimates even under highly rare traffic states, exhibiting superior worst-case robustness.
@@ A.1 Risk-Sensitive RL: The exponential barrier
A.2 RE-SAC: A heuristic argument for polynomial scaling
Why might RE-SAC avoid the exponential barrier? The key observation is that its penalty terms are deterministic functions of the network weights (for $\kappa$) or the frozen target ensemble (for $\Gamma_{epi}$), rather than statistics of the return distribution.
Directly Optimizing $\mathbb{E}[\exp(\beta R)]$ (Risk-Sensitive) requires accurate estimation of the tail distribution, which is statistically expensive (hence $\exp(H)$). However, RE-SAC leverages the structural correspondence shown by Xu et al. (Appendix B), which maps IPM-based robustness to weight-norm regularization.
@@ A.2 RE-SAC: A heuristic argument for polynomial scaling
This transformation converts the “Hard” Risk-Sensitive problem into a “Standard” Risk-Neutral MDP with augmented reward $\tilde{R}(s,a)=R(s,a)-\lambda\|W\|^{2}$.
Informal argument. If the penalties are treated as fixed (which they are per Bellman backup; see §E), the RE-SAC operator acts on a standard shifted-reward MDP whose tabular regret scales as $\tilde{O}(\sqrt{H^{3}S^{2}AT})$, polynomial in the horizon HH—in contrast to the exponential $\exp(|\beta|H)$ dependence of risk-sensitive approaches.
Important caveat. This comparison is an informal motivation, not a formal regret theorem for RE-SAC, for two reasons:
  1. 1.
    The penalties $\kappa$ and $\Gamma_{epi}$ change across training iterations as the network weights and target ensemble evolve. A rigorous regret analysis would need to account for this non-stationarity (cf. the per-step discussion in §E.1).
  2. 2.
    The shifted reward $\tilde{R}$ may have a different range than $R$, affecting the constants in the bound.
The conceptual takeaway is that by replacing tail-distribution estimation (which requires exponential samples) with a deterministic structural penalty (which requires no additional samples), RE-SAC sidesteps the mechanism that causes the exponential barrier.
Appendix B The correspondence chain: From risk to regularization